OpenStudy (dls):

If the radius of the earth shrinks by 5%,mass remaining the same,then how much value of acc. due to gravity change?

OpenStudy (anonymous):

radius shrinks by 5% means.. the radius becomes 95/100(Re - current earths radius).. So can you solve now?

OpenStudy (dls):

i know that much i did that too didnt get the answer though

OpenStudy (anonymous):

wait... so you tell me now.. once you substitue this new value of radius.. what is the ratio of the new value of g to the present value of g?

OpenStudy (dls):

i have to give the answer is "%" change

OpenStudy (anonymous):

Ok we ll work towards it don't worry.. !

OpenStudy (anonymous):

answer my question first :P

OpenStudy (dls):

6400/6080 ?

OpenStudy (anonymous):

:O.. no.. how did you get that? :-/

OpenStudy (anonymous):

listen .. use the equation of g..

OpenStudy (dls):

\[\frac{95}{100} \times 6400\]

OpenStudy (dls):

New one^^

OpenStudy (anonymous):

yes.. but i want the value of g.. not the radius!

OpenStudy (dls):

\[g= \frac{GM}{R ^{2}}\]

OpenStudy (dls):

R'=95/100R

OpenStudy (anonymous):

yes exactly.. now put the new value of R .. and get the ratio!!

OpenStudy (anonymous):

correct.. now once you put new value.. you ll get a new g.. lets call that g prime... so find out what is the ratio of gprime to g?

OpenStudy (dls):

20GM/19R is the new one? :S

OpenStudy (anonymous):

if you plug in the values.. you should get g (prime)/ g = 1.108

OpenStudy (dls):

yeah..okay

OpenStudy (dls):

so 11%?

OpenStudy (anonymous):

10.8 percent..

OpenStudy (dls):

11.08%?

OpenStudy (anonymous):

no.. see.. the increase is .108 ( 1.108-1) ... hence percent is 10.8!

OpenStudy (dls):

okayy

OpenStudy (dls):

answer is 10% though!

OpenStudy (anonymous):

wait . lemme see if i did some calc. mistake!

OpenStudy (dls):

the solution is solved using differentiation n stuff

OpenStudy (anonymous):

!? :-/ why differentiation? :O :O

OpenStudy (dls):

DK

OpenStudy (dls):

\[\frac{dg}{dR} = \frac{d}{dR} ( \frac {GM}{R ^{2}} ) = GM \frac{d}{dR} (R^{-2}) \]

OpenStudy (dls):

Solving this we got, \[\frac{-2g}{r}\]

OpenStudy (anonymous):

ohh.. like that.. ! thats not really required!!

OpenStudy (dls):

dg/g=-2dR/R

OpenStudy (dls):

:O

OpenStudy (dls):

at last we get 0.1*100=10%!!

OpenStudy (anonymous):

i mean it can be done in the normal method i told :P.. but this method holds too.. i dunno why we getting answer difference :D.. maybe something is wrong!! lemme recheck and get back

OpenStudy (anonymous):

i find no flaw in calc.. i dunno really what went wrong :-/

OpenStudy (vincent-lyon.fr):

There is a difference, because the differential method gives only a fist-order approximation of the result. The real answer is the one in which you work out the new value of g. If the shrinking had been 20%, the differential method would have lead to a very wrong value. On the other hand, if the shrinking were 0.1%, then your calculator would probably have problems coping with two big numbers almost identical to subtract, whereas the differential method would give a reliable answer.

OpenStudy (anonymous):

ow.. thaks vincent for clearing that up! But why does the differential method give only an approximation!?... dg/dr, means change in the value of g for a change in value of r.. so why is it an approximation?

OpenStudy (vincent-lyon.fr):

 My dr should have been negative to match your problem. If dr is too big, the first order differential can lead you (by following the tangent and not the curve) to a value far from the real one.